OpenStudy (anonymous):

SOLVE 4x - 8 = 2(2x - 3)

OpenStudy (cj49):

4x-8=4x-6 x=2

OpenStudy (anonymous):

So You have 4x - 8 = 2(2x - 3) Simplify : 4 (x - 2) = 4x - 6 4 (x - 2) = 2 (2x - 3) So your answer would be x = 2 Hope this helped

OpenStudy (anonymous):

kind of but that's not any of my choices so i'm confused